OpenStudy (anonymous):

Solve. 12.6-0.4h+2.5-0.5=15-0.9h-7.1

OpenStudy (anonymous):

12.6+2.5-.5=14.6 15-7.1=7.9 so simply you get 14.6-.4h=7.9-.9h so now you need to get like terms together so bring the number with the variable h to the left of equal sign and whole # 14.6 to right: add .9h to both sides and substract 14.6 from both side of equal sign .9h-.4h=-14.6+7.9 .5h=-6.7 divide by .5 h=-6.7/.5

OpenStudy (anonymous):

Thank you @kmalone99